\left(x-y\right)dx+\left(3x+y\right)dy=0


1
Expert's answer
2021-10-17T17:46:49-0400

Let us solve the differential equation "\left(x-y\right)dx+\left(3x+y\right)dy=0." For this let us use the transformation "y=ux." Then "dy=udx+xdu," and we get the equation "\left(x-ux\right)dx+\left(3x+ux\right)(udx+xdu)=0." Since "x=0" is not a solution, let us divide both parts by "x." Then we get "\left(1-u\right)dx+\left(3+u\right)(udx+xdu)=0," which is equivalent to "\left(1+2u+u^2\right)dx+\left(3+u\right)xdu=0." It follows that "\frac{dx}x+\frac{\left(3+u\right)du}{(1+u)^2}=0," and hence "\int\frac{dx}x+\int \frac{\left(1+u+2\right)du}{(1+u)^2}=C." We conclude that "\ln|x|+\int \frac{du}{1+u}+2\int \frac{du}{(1+u)^2}=C," and hence "\ln|x|+\ln|1+u|-2\frac{1}{1+u}=C."

We conclude that "\ln|x|+\ln|1+\frac{y}x|-\frac{2x}{x+y}=C" is the general solution of the equation "\left(x-y\right)dx+\left(3x+y\right)dy=0."